OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(–4). A. -5 B. 14 C.12 D. -10

OpenStudy (ipwnbunnies):

Plug in -4 whereever you see x.

OpenStudy (ipwnbunnies):

The biggest trick here is to get the signs right.What'd you get?

OpenStudy (anonymous):

i got 14 -4=2-3(-4) 3*(-4)=12 2-(-12)=14

OpenStudy (ipwnbunnies):

Good job! And it'll be f(-4) to the left of the equation.

OpenStudy (anonymous):

kk thxxx